Jack takes 7.5 hours to make a trip overseas in a prop plane. Later he discovers that he could have taken a jet and saved 2.5 hours of flying time. Find the speed of the jet if the jet travels 225 MPH faster than the prop plane. -Thank you soo much!
Answer by scott8148(6628)

d = r * t ___ the distance is the same for both planes

(j - 225) * 7.5 = j * (7.5 - 2.5)

7.5j - (7.5 * 225) = 5j

2.5j = 7.5 * 225 ___ j = 675